# Need for zkSpeed: Accelerating HyperPlonk for Zero-Knowledge Proofs

### v1

@@ 4.3. Multifunction Tree Unit
Many HyperPlonk functions exhibit binary-tree compute patterns, including building MLE, MLE evaluation, and constructing Product MLE (). zkSpeed supports these in hardware with our multifunction tree unit, specially designed to handle these compute patterns effectively. Building MLE is a function used in ZeroCheck and Opencheck steps. It constructs a table with entries from random values to , where represents problem size. The computation is divided into layers to reduce the number of modular multiplications from times to . The multiplier tree is used for batch inversion during Fraction MLE generation, where it efficiently calculates the product for an inversion batch size, . Similarly, the MLE evaluation in the Batch Evaluation operates like a multiplier tree but includes additional modular additions in each operation. The Product MLE generation in the Wiring Identity outputs all layer results. The functions’ dataflow are presented in Figure 2.
The original (CPU) HyperPlonk implementation uses breadth-first (level-order) tree traversal (BFS). This is inappropriate for hardware acceleration as it puts increased pressure on SRAM capacity and off-chip bandwidth. For example, a problem size of requires up to intermediate elements in a layer, each 255 bits wide, which would require 128MB for the intermediates alone. We propose to use depth-first traversal (DFS) to address these challenges. This approach reuses and consumes intermediate results as they are computed, reducing the up to 255MB intermediates that must be stored on-chip or spilled to DRAM for problem size .
As shown in Figure 2, for Build MLE, if the PEs produce two outputs per cycle at the last layer, other PEs will keep generating one element each cycle in the previous layer, and one element every two cycles of the first layer. Outputs are produced continuously in the last layer, and intermediate results like are discarded after use. The multiplier tree (in FracMLE) and MLE evaluation also benefit from DFS. For instance, as long as the unit reads two MLE table entries each cycle, it can keep generating one item of the first layer per cycle, one item every two cycles of the second layer, and so on to sustain computation similarly to a pipeline. For ProdMLE, the design outputs all layers. Since we generate an item of the second layer during the first layer, we label each item’s layer index to store results correctly.
@@
Figure 5 shows how the multifunction tree unit works. Each PE includes a modular multiplier and modular adder. (E.g., for Build MLE, and only require one modular multiplication since is computed as ). The hardware supports three processing modes of Figure 2. For the Inverse Tree (e.g., MLE Evaluation), the unit accepts inputs in parallel and reduces tree-level partials via a matching hardware dataflow; in Figure 5, , and the data flows from left to right. If the tree has more levels than hardware supports (three levels in the example; a problem size workload has 20 levels), the remaining levels are scheduled and processed via the accumulator. Outputs from the hardware tree (Level 3 in Figure 5) are pushed in an accumulator-local register file and once operands are ready, popped to the accumulator PE. The bottom of Figure 5 shows its schedule. Initially, there are gaps as the accumulator must wait two cycles for each input pair. However, once multiple levels are processed, the gaps are filled; this can be seen in cycle 44, where layers 4 and 5 are processed by the PE at the same time (i.e., both in the pipeline). Thus the PEs in the tree have high utilization: for a workload, they are over 99% utilized during the whole computation. The red arrows coming out of each PE show how computations are sampled and output to support ProdMLE. Forward Tree (Build MLE) support is shown in blue, and data flows from right to left. Each PE takes the previous level and a challenge as inputs and generates two outputs that are fed to the next level. Similarly, if the tree has more levels, the accumulator PE is scheduled to generate (roughly) one output per cycle to feed the rest of the PEs, which correspond to the last 3 levels, outputting 8 results in the end. Switching is supported by muxes at the PE inputs and configuring them.
@@
The advantages of DFS are noticeable: it eliminates the need to store entire intermediate layers, making it practical for large problem sizes, and provides the ability to rate-match with upstream or downstream units and maintain throughput. By adjusting the number of PEs, the unit can handle varying input and output rates, forming a full pipeline with other units. The ability to reuse across multi-functions eliminates the need to allocate multiple dedicated units, saving 41.6% area across global Pareto design points in Section 7. The HyperPlonk code uses BFS. BFS has greater dependence distance; we tile DFS and schedule work to avoid dependence stalls. Executing nodes already stresses CPU’s limited compute resources, and we expect DFS to have little performance impact in software.
4.4. Construct N&D and Fraction MLE
@@
The construct N&D stage generates the and MLEs discussed in Section 3.3.3. Elements of six intermediate MLEs, , and , are computed in parallel from modular additions and multiplications of the witness and wiring permutation MLEs stored in on-chip SRAM, and two random challenges from SHA3. These intermediate MLEs are written off-chip for the subsequent PermutationCheck. The intermediate MLEs are multiplied to obtain the and MLE (e.g., ) elements, and fed to the FracMLE unit.
@@
Fraction MLE, , requires computing the modular inverse of every element of the Denominator MLE table (), and multiplying each inverted element with the corresponding element of the Numerator MLE table (). Given , modular inversion outputs a , such that . We use the constant time Binary Extended Euclidean Algorithm (BEEA) (Pornin, 2020) for this operation to make it data-oblivious and ensure constant scheduling. The algorithm requires iterations of a loop consisting of shifts and subtractions, where is the number of input bits (), resulting in a 509-cycle latency. A data-dependent implementation is faster for smaller input values, but since we are computing on random inputs (derived from SHA3 hashes), the average latency of this implementation (2 cycles) is only 1 better. In exchange for this negligible latency overhead, our constant-time implementation reduces design complexity as this ensures elements of (and thus FracMLE) are generated in-order if we execute multiple inverses in parallel (as we will discuss shortly). Parallel execution of the data-dependent algorithm may output elements out-of-order and would require buffering or stalling to resolve.
This is an expensive operation for HyperPlonk as 255-bit elements are used, and each element of an MLE must be inverted, so we heavily optimize it. We leverage Montgomery batching (Montgomery, 1987), which allows us to compute the inverse of multiple elements at the cost of one inversion and additional multiplications, to amortize the cost of one inversion across multiple elements and improve per-inversion throughput. The common optimized approach to batching stores the partial products generated while computing the overall product using sequential multiplications. After inversion, the inverted output is sequentially multiplied by batch elements to isolate inverted batch elements. The sequential multiplications, combined with the latency of modular multiplication, can limit performance for hardware implementations. We address these limitations with two modifications. First, we use a multiplier tree, detailed in Section 4.3; this significantly reduces latency and improves scaling for large values of , from to multiplications. Second, we pre-compute the sequential multiplications needed to obtain the partial products in parallel to the modular inverse operation. This hides the latency of sequential multiplications and maximizes the work done during the long latency of the modular inverse.
@@
Figure 6 illustrates the architecture of our modular inverse unit using batching. We use multiple batched inversion units in round-robin fashion to completely mask long inversion latencies and enable the FracMLE unit to accept one input and generate one output per cycle, behaving as a pipeline with depth . We achieve this by using enough batched inverse units to mask the latency of one batch inversion. The multiplier tree and one multiplier are reused across all units to compute their inputs and produce the individual elements of , respectively. One batch inversion latency is the maximum of the parallel partial product latency and the combined multiplier tree & modular inverse latencies. The former scales while the latter scales .
We frame choosing the ideal as an optimization problem to minimize the imbalance between the two latencies. The left y-axis of Figure 7 plots the latency imbalance as a function of . The initial latency imbalance is due to the high constant latency of modular inverse, which decreases as the partial product latency increases – reaching a minimum at (64) – then increases as the partial product latency overtakes the multiplier tree and modular inverse latency. Another consideration when choosing is the area cost due to the multiplier and storage overhead. The right y-axis of Figure 7 plots the total area of the Modular Inverse unit for different batch sizes, including the multiplier tree, partial products SRAM, batch element registers, and multipliers. The minimum area is also seen at 64 because more inverse units are needed to mask the latency of one inversion at smaller batch sizes. For example, requires 256 batched inverse units while only 12 units are required at . Also, starting at , we can reuse the multiplier tree across all units since the tree can compute a batch’s product before the next batch is ready thanks to its log scaling. The total area increases past because the latency of the multiplications for the partial products overtakes the latency of modular inversion (hence latency imbalance also starts to increase), and the number of batched inverse units no longer decreases while the area overhead of batching (e.g., SRAM) continues to increase without additional benefit. Using these metrics, we get a batch size of 64.
@@ 4.5. MLE Combine Unit
The MLE Combine Unit is used in the Polynomial Opening step. As mentioned in Section 3.3.5, there are several linear combinations of MLEs that are performed before OpenCheck and before the MSMs. These operations are straightforward, and use a combination of MLEs stored in on-chip SRAM and off-chip memory to construct the MLEs used in subsequent steps. Because Opencheck happens in series with the MSMs, the respective MLE Combine operations also happen in sequence. Consequently, we can share resources between these two operations. For the design point we highlight in Table 5, without sharing, we would require 122 modular multipliers. With sharing, we require only 72, representing a 41% area savings.
@@ 4.6. On-Chip MLE SRAM
The MLE table size scales with the number of gates, i.e., for a problem size of gates, each MLE table has entries. In practice, input MLEs are sparse. Control MLEs are all binary, and are roughly 90% 1s and 0s and 10% full bit-width. These MLEs get reused throughout the protocol, so we store them on-chip in Global SRAM. We compress the tables, packing together control MLEs and using address translation units to perform lookups to either binary or 255-bit data. These compression strategies save to on MLE storage across problem sizes.
@@
Figure 8 provides an overview of the HyperPlonk protocol and zkSpeed architecture. HyperPlonk is expressed as a series of five steps, with the final steps (4A and 4B) executed in parallel. Colors indicate the mapping of protocol steps (Section 3.3) to the accelerator units (Section 4) they run on. The zkSpeed architecture is streaming in nature and captures on-chip data reuse when feasible via explicitly managed scratchpad memories. The architecture has four major components: accelerator units (Section 4), local and global SRAM, a multi-channel shared bus, and HBM interface. HBM is needed to feed the chip with high bandwidths needed by HyperPlonk, and we conduct bandwidth sensitivity studies in Section 7.
zkSpeed uses a shared bus rather than a crossbar or NoC. This design choice was made after rigorously analyzing the HyperPlonk dataflow. The dataflow of each HyperPlonk protocol step is shown in Figure 8(C). Colors indicate the hardware module each kernel executes on and annotated wires show data movement. We observed that at any given time, only 1-2 zkSpeed units typically communicate, and at most 4 independent bus channels are needed to avoid stalling – this is during Wire Identity where Construct N&D sends results to FracMLE, FracMLE simultaneously feeds ProdMLE and MSM units, and ProdMLE streams to MSM. Units overlap computation with each other, e.g., enabling MSM to start processing partial outputs from FracMLE, effectively masking latency. Without bus stalls, we are able to rate match each accelerator unit to pipeline across modules when possible, further improving performance.
@@
Dataflow. HyperPlonk is data oblivious at the stage granularity. This allows it to statically schedule computations and manage units, SRAM, and the buses via a simple controller. The Batch Evaluation and Polynomial Opening steps run in parallel because the same MLEs used prior to Opencheck also undergo MLE evaluations. Both steps require simultaneous access to 13 MLE tables. With our on-chip MLE storage scheme, only two of these MLEs are stored off-chip, and . This saves bandwidth by 84%. By synchronizing the off-chip accesses between the two protocol steps, zkSpeed further halves the bandwidth pressure. We provision six hardware units for computing Batch Evaluations, since this step is not on the critical path, and the remaining MLE evaluations only rely on MLEs already stored on-chip. Thus we eliminate batch evaluations from the critical path without incurring additional bandwidth costs.
@@
Hyperplonk Trends and Outlook. Zero-knowledge protocols are still continuously evolving and improving in several aspects. However, zkSpeed, with its focus on Hyperplonk still has significant stability. Firstly, Hyperplonk has seen adoption both in industry implementations (Zhang et al., [n. d.]; Diamond and Posen, 2023) and academic research (Kohrita and Towa, 2024; Diamond and Posen, 2023, 2024). In addition, our design is modular, and the key components of Hyperplonk SumChecks, MSMs, and MLEs, are present in essentially any modern SNARK protocol (Thaler, 2013; Setty, 2020; Groth, 2016; Setty et al., 2024; Kothapalli et al., 2022; Bünz and Chen, 2023). Therefore, zkSpeed can also be targeted to new protocols so long as they comprise the fabricated modules, including proof composition methods (Valiant, 2008) that seek to compose protocols like Orion and Hyperplonk.
@@ 6.1. Performance Modeling
SumCheck has a fixed dataflow and is data oblivious. Thus, we model performance using analytical models. Modules that feed directly to the SumCheck units are handled in a data-oblivious manner, so they are also modeled analytically since we assume rate-matching. For the MSM, we use a cycle-accurate simulator to model performance. Each unit within zkSpeed is modularized. To understand full-chip performance, we conduct a design-space exploration of all combinations of design parameters detailed in Table 2 and then analyze the pareto-optimal space to pick a suitable configuration for profiling runtimes. We also construct power traces to estimate average power for the full-chip architecture. We use Catapult HLS 2023 to generate the RTL for Montgomery multipliers (as done in prior work (Daftardar et al., 2024)), the fully-pipelined, unified SumCheck unit (to handle ZeroCheck, PermutationCheck, and OpenCheck), and the fully-pipelined PADD unit. Consistent with HyperPlonk, we use the BLS12-381 elliptical curve, where all MLE datatypes (e.g., in the SumCheck unit) are 255b, and all elliptic curve points (e.g., in the PADD) are 381b. Using Design Compiler with TSMC 22nm, we find the critical path in our design is the 381b PADD unit at 1.05ns. We use Synopsys 22nm Memory Compiler to generate SRAM estimates. For SHA3, we use the publicly available IP block from OpenCores (OpenCores, 2013). We scale down to 7nm using scale factors of 3.6 for area, 3.3 for power, and 1.7 for delay (as in prior work (Daftardar et al., 2024)), and clock all zkSpeed accelerators at 1 GHz.

### v2

@@ 4.3. Multifunction Tree Unit
4.3.1. Tree-like Computation in HyperPlonk
Many HyperPlonk functions exhibit binary-tree compute patterns, including Build MLE, MLE Evaluate, and constructing Product MLE (). zkSpeed supports these in hardware with our Multifunction Tree Unit (MTU), specially designed to handle these compute patterns effectively. Build MLE is a function used in ZeroCheck and OpenCheck steps. It constructs a table with entries from random values to , where represents problem size. The computation is divided into layers (as a binary tree) to reduce the number of modular multiplications from to . The multiplier tree is used for batch inversion during Fraction MLE generation (see Section 4.4), where it efficiently calculates the product for an inversion batch size, . Similarly, the MLE Evaluate in the Batch Evaluation step operates like a multiplier tree but includes additional modular additions in each operation. The Product MLE generation in the Wiring Identity outputs all layer results. The functions’ dataflows are presented in Figure 3.
4.3.2. zkSpeed’s Tree Approach
The original (CPU) HyperPlonk implementation uses breadth-first (level-order) traversal (BFS). This is inappropriate for hardware acceleration as it puts increased pressure on SRAM capacity and off-chip bandwidth. For example, a problem size of requires up to intermediate elements in a level, each 255 bits wide, which would require 128MB for the intermediates alone. We propose a hybrid strategy that applies depth-first traversal (DFS) to the upper levels of the binary tree and switches to BFS at lower levels. This approach leverages efficient memory reuse in the upper levels through DFS, while enabling parallelism in the lower levels via BFS. It reuses and consumes intermediate results as they are produced, reducing the amount of data that must be stored on-chip or spilled to DRAM.
@@ 4.3.3. Multifunction Tree Architecture
We further analyze this tree architecture in (Mo et al., 2025a). Figure 6 shows an example of how the MTU works. Each PE includes a modular multiplier and modular adder444For Build MLE, and only require one modular multiplication since is computed as . The hardware supports three processing modes of Figure 3. For the Inverse Tree (e.g., MLE Evaluate), the unit accepts inputs in parallel and reduces tree-level partials via a matching hardware dataflow; in Figure 6, , and the data flows from left to right. If the tree has more levels than hardware supports (three levels in the example; a problem size workload has 20 levels), the remaining levels are scheduled and processed in DFS via the accumulator. Outputs from the hardware tree (Level 3 in Figure 6) are pushed in an accumulator-local register file and once operands are ready, popped to the accumulator PE. The bottom of Figure 6 shows its schedule. Initially, there are gaps as the accumulator must wait two cycles for each input pair. However, once multiple levels are processed, the gaps are filled; this can be seen in cycle 44, where levels 4 and 5 are processed by the PE at the same time (i.e., both in the pipeline). Thus the PEs in the tree have high utilization: for a workload, they are over 99% utilized during the whole computation. The red arrows coming out of each PE show how computations are sampled and output to create the Product MLE. The Forward Tree (Build MLE) support is shown in blue, and data flows from right to left. Each PE takes the previous level and a challenge as inputs and generates two outputs that are fed to the next level. Similarly, if the tree has more levels, the accumulator PE is scheduled to generate (roughly) one output per cycle to feed the rest of the PEs, corresponding to the last 3 levels, outputting 8 results in the end. Switching is supported by muxes at PE inputs.
@@
The advantages of this hybrid traversal (Mo et al., 2025a) are noticeable: it eliminates the need to store entire intermediate levels, making it practical for large problem sizes, and provides the ability to rate-match with upstream or downstream units and maintain throughput. By adjusting the number of PEs, the unit can handle varying input and output rates, forming a full pipeline with other units. The ability to reuse across multi-functions eliminates the need to allocate multiple dedicated units, saving 41.6% area across global Pareto design points in Section 7. As HyperPlonk’s code uses BFS, it experiences greater dependence distances. Our traversal mitigates this by scheduling work to avoid dependence stalls. Executing nodes already stresses CPU’s limited compute resources, so we expect hybrid traversal to have little performance impact in software.
4.4. Construct N&D and FracMLE
4.4.1. Creating with Constant-Time Inversion
@@
The Construct N&D stage generates the and MLEs discussed in Section 3.3.3. Elements of six intermediate MLEs, , and , are computed in parallel from modular additions and multiplications of the witness () and wiring permutation () MLEs stored in on-chip SRAM, and two challenges from SHA3. These intermediate MLEs are written off-chip for the subsequent PermCheck and multiplied to obtain the and (e.g., ) elements, and fed to the FracMLE unit.
@@
The Fraction MLE, , requires computing the modular inverse of every element of the Denominator MLE (), and multiplying each inverted element with the corresponding element of the Numerator MLE (). Given , modular inversion outputs a , such that . We use the constant time Binary Extended Euclidean Algorithm (BEEA) (Pornin, 2020) for this operation to make it data-oblivious and ensure constant scheduling. The algorithm requires iterations of a loop consisting of shifts and subtractions, where is the number of input bits (), resulting in a 509-cycle latency. A data-dependent implementation is faster for smaller input values, but since we are computing on random inputs (derived from SHA3 hashes), the average latency of such an implementation (2 cycles) is only 1 better. In exchange for this negligible latency overhead, our constant-time implementation reduces design complexity as this ensures elements of (and thus ) are generated in-order if we execute multiple inverses in parallel (as we will discuss shortly). Parallel execution of the data-dependent algorithm may output elements out-of-order and would require buffering or stalling to resolve.
4.4.2. Batching for Modular Inversion
Computing modular inverses for each 255-bit MLE element is expensive, so we heavily optimize it. We leverage Montgomery batching (Montgomery, 1987) to amortize the cost of one inversion across multiple elements and improve per-inversion throughput. In the standard approach (also implemented in (Chen et al., 2022)), to compute the inverses of a batch of elements (e.g., ), partial products are computed sequentially (e.g., , , , ) and the final product is inverted. This single inverse is then propagated backward to recover the individual inverses (, etc.) through additional sequential multiplications. While this reduces the number of inversions, the sequential multiplications and high modular multiplication latency can limit hardware performance.
We address these limitations with two modifications. First, we use a multiplier tree, detailed in Section 4.3; this significantly reduces latency and improves scaling for large values of the batch size, which we denote , from to multiplications. Second, we overlap the sequential multiplications for partial products with modular inversion to mask long multiplication latencies.
4.4.3. Batch Inversion Architecture
@@
Figure 7 illustrates the architecture of our modular inverse unit using batching. We use multiple batched inversion units in round-robin fashion to completely mask long inversion latencies and enable the FracMLE unit to accept one input and generate one output per cycle, behaving as a pipeline with depth where is the number of FracMLE units. We achieve this by using enough batched inverse units to mask the latency of one batch inversion. The multiplier tree and one multiplier are reused across all units to compute their inputs and produce the individual elements of , respectively. One batch inversion latency is the maximum of the parallel partial product latency and the combined multiplier tree and modular inverse latencies. The former scales while the latter scales .
4.4.4. Optimizing for Batch Size
We frame the choice of batch size as an optimization problem to minimize latency imbalance (between partial products and the multiplier tree and inversion) and area. As shown in Figure 8, the left y-axis plots latency imbalance, which is initially high due to the fixed cost of modular inversion. It decreases as partial product latency rises, reaching a minimum at , then increases again as partial-product latency begins to dominate. The right y-axis shows total area (including multiplier trees and partial products SRAM), which also reaches a minimum at due to fewer required inverse units (e.g., 256 units for vs. only 12 for ). Additionally, starting at , we can reuse the multiplier tree across all units since its latency allows it to complete one batch before the next arrives. Beyond , partial-product latency surpasses modular inversion, and SRAM overheads continue to grow without reducing the number of inverse units—diminishing returns in both performance and area. Based on these tradeoffs, we select as the optimal batch size.
@@ 4.5. MLE Combine Unit
The MLE Combine Unit is used in the Polynomial Opening step. As mentioned in Section 3.3.5, there are several linear combinations of MLEs that are performed before OpenCheck and before the MSMs. These operations are straightforward, and use a combination of MLEs stored in on-chip SRAM and off-chip memory to construct the MLEs used in subsequent steps. Because OpenCheck happens in series with the MSMs, the respective MLE Combine operations also happen in sequence. Consequently, we can share resources between these two operations. For the design point we highlight in Table 5, without sharing, we would require 122 modular multipliers. With sharing, we require only 72, representing 41% area savings.
@@ 4.6. On-Chip MLE SRAM
The MLE table size scales with the number of gates, i.e., for a problem size of gates, each MLE table has entries. In practice, input MLEs are sparse. Control MLEs are all binary, and are roughly 90% 1s and 0s and 10% full bit-width. These MLEs get reused throughout the protocol, so we store them on-chip in global SRAM. We compress the tables, packing together control MLEs and using address translation units to perform lookups to either binary or 255-bit data. These compression strategies save to on MLE storage across problem sizes. This also significantly reduces off-chip accesses; for example, in the Polynomial Opening step, only 2 of 13 MLE tables ( and ) are stored off-chip, cutting bandwidth usage by 84%.
@@
Figure 2 provides an overview of the HyperPlonk protocol and zkSpeed architecture. Colors indicate the mapping of protocol steps (Section 3.3) to the accelerator units (Section 4) they run on. The zkSpeed architecture is streaming in nature and captures on-chip data reuse when feasible via explicitly managed scratchpad memories. The architecture has four major components: accelerator units, local and global SRAM, a multi-channel shared bus, and HBM interface. HBM is needed to feed the chip with high bandwidths needed by HyperPlonk, and we conduct bandwidth sensitivity studies in Section 7.
Dataflow. HyperPlonk is data oblivious at the stage granularity. This allows zkSpeed to statically schedule computations and manage units, SRAM, and the buses via a simple controller. Further, zkSpeed uses a shared bus rather than a crossbar or NoC. This design choice was made after rigorously analyzing the HyperPlonk dataflow (see Figure 2C). We observed that at any given time, only 1-2 zkSpeed units typically communicate, and at most 4 independent bus channels are needed to avoid stalling – this is during Wire Identity where Construct N&D sends results to FracMLE, FracMLE simultaneously feeds ProdMLE and MSM units, and ProdMLE streams to MSM. Units overlap computation with each other, e.g., enabling MSM to start processing partial outputs from FracMLE, effectively masking latency. Without bus stalls, we are able to rate match each accelerator unit to pipeline across modules when possible, further improving performance.
@@
HyperPlonk Trends and Outlook. Zero-Knowledge protocols are still continuously evolving and improving in several aspects. However, zkSpeed, with its focus on HyperPlonk, still has significant stability. Firstly, HyperPlonk has seen adoption both in industry implementations (Zhang et al., 2022; Diamond and Posen, 2023) and academic research (Kohrita and Towa, 2024; Diamond and Posen, 2023, 2024). In addition, our design is modular, and the key components of HyperPlonk SumChecks, MSMs, and MLEs, are present in essentially any modern SNARK protocol (Thaler, 2013; Setty, 2020; Groth, 2016; Setty et al., 2024; Kothapalli et al., 2022; Bünz and Chen, 2023). Therefore, zkSpeed can also be targeted to new protocols so long as they comprise the fabricated modules, including proof composition methods (Valiant, 2008) that seek to compose protocols like Orion and HyperPlonk.
@@ 6.1. Performance Modeling
SumCheck has a fixed, data-oblivious dataflow, which allows us to model its performance analytically. Modules that feed into the SumCheck units are also data-oblivious and are modeled similarly, accounting for rate-matching assumptions. For the MSM, we use a cycle-accurate simulator to model performance. Each unit within zkSpeed is modularized. To understand full-chip performance, we conduct a design space exploration of all combinations of design parameters detailed in Table 2 and then analyze the pareto-optimal space to pick a suitable configuration for profiling runtimes. We also construct power traces to estimate average power for the full-chip architecture. We use Catapult HLS 2023 to generate the RTL for Montgomery multipliers (as done in prior work (Daftardar et al., 2024)), the fully-pipelined, unified SumCheck unit (to handle ZeroCheck, PermCheck, and OpenCheck), and the fully-pipelined PADD unit. Consistent with HyperPlonk, we use the BLS12-381 elliptical curve, where all MLE datatypes (e.g., in the SumCheck unit) are 255-bit, and all elliptic curve points (e.g., in the PADD) are 381-bit. Using Design Compiler with TSMC 22nm, we find the critical path in our design is the 381-bit PADD unit at 1.05ns. We use Synopsys 22nm Memory Compiler to generate SRAM estimates. For SHA3, we use the IP block from OpenCores (OpenCores, 2013). We scale down to 7nm using scale factors of 3.6 for area, 3.3 for power, and 1.7 for delay (as in prior work (Daftardar et al., 2024)), and clock all zkSpeed accelerators at 1 GHz.